OpenStudy (anonymous):

what is the standard deviation of the data set given below 2 3 6 9 10 A: 10 B: square root of 10 C: square root of 12.5 D: square root of 3.5 E: 3.5 F: 12.5

OpenStudy (kirbykirby):

Standard deviation: \[\Large = \sqrt{\frac{1}{n-1}\sum_{i=1}^n(x_i-\bar{x})^2} \]

OpenStudy (anonymous):

can you tell me what to plug in where because i dont understand the formula

OpenStudy (kirbykirby):

\(n\) is the sample size (the # of data points you have), which is 5 \(x_i\) is referring to each data point \(\bar{x}\) is the mean of your data so \(x_i - \bar{x}\) means to take a data point and subtract it from the mean \[\sum_{i=1}^n (x_i-\bar{x})\] means take the 1st point and subtract it with the mean, then add that to (the 2nd point subtracted to the mean) and keep doing that until the fifth point. \[\sum_{i=1}^n (x_i-\bar{x})^2\] is similar to what I describe above, except when you take the differences, you must square that result. In your example: The mean is \[ \frac{2+3+6+9+10}{5}=6\] So the formula, once plugged in, should like like: \[ \sqrt{\frac{1}{5-1}\left( (2-6)^2+(3-6)^2+(6-6)^2+(9-6)^2+(10-6)^2 \right)}\]

OpenStudy (anonymous):

Thank you so much! it helped a ton!

OpenStudy (kirbykirby):

awesome :)

OpenStudy (campbell_st):

that's is the sample standard deviation, I hope you don't need to population standard deviation. I always thought the population was the default

OpenStudy (kirbykirby):

hmm, I was always under the assumption that the sample one was the default. Population data are rare to come by. But I guess it's still a valid point. Hopefully your teacher told you what to assume. If it's a population standard deviation, you should use \(\Large \frac{1}{5}\) instead of \(\Large \frac{1}{5-1}\)

OpenStudy (anonymous):

square root 12.5